What is the empirical formula for a compound which contains 0.0134g of iron, 0.00769g of sulfur, and 0.0115g of oxygen
1
Expert's answer
2018-02-01T05:05:19-0500
m (Fe)=0.0134 g
m (S)=0.00769 g
m (O)=0.0115 g
Fe x S y O z=?
We find the amount of the substance of each element:
. n (Fe)= m(Fe)/Ar (Fe)=0.0134/56=0.00024 mol
. n (S)= m (S)/Ar(S)=0.00769/32=0.00024 mol
.n(O) =m(O)/Ar(O)=0.0115/16=0.00072 mol
. n(Fe) : n(S): n(O)=0.00024:0.00024:0.00072= 1:1:3
The formula of the compound FeSO3 –iron sulphite